Two spheres of same material and radii 5m and 2m are at temperature 200K and 2500K respectively.The ratio of energies radiated by them per second is
1
Expert's answer
2021-08-25T18:19:08-0400

Gives

"r_1=5m\\r_2=2m\\T_1=200K\\T_2=2500K"

We know that stefan's law

"E=\sigma AT^4"

"\frac{E_1}{E_2}=\frac{A_1T_1^4}{A_2T_2^4}"

"\frac{E_1}{E_2}=\frac{r_1^2T_1^4}{r_2T_2^4}"

"\frac{r_1}{r_2}=\frac{5}{2}"

"\frac{E_1}{E_2}=\frac{25}{4}\times(\frac{200}{2500})^4=\frac{4}{15625}"